geerky42 (geerky42):

How many numbers in the form \(a^4\), where \(a \in \mathbb{Z}^+\) divide \(3! \times 4! \times 7!\) ?

OpenStudy (anonymous):

i don't think there are too many since you need primes to the power of 4

OpenStudy (anonymous):

included in \(3!4!7!\) is \(2^7\) and \(3^4\) all other primes are to lower powers

OpenStudy (anonymous):

i am not certain but on the basis of prime factorization i only see \[2^4,3^4,(2\times 3)^4\]

OpenStudy (anonymous):

oops i miscounted!! it is \(2^8\) and \(3^4\)

OpenStudy (anonymous):

so maybe there are 4 all together, \[2^4, 3^4, (2^2)^4,(2\times 3)^4, (2^2\times 3)^4\]

OpenStudy (anonymous):

well that is actually 5, not 4